# Moving Out: Physically-grounded Human-AI Collaboration

### v1

@@ C.1 Hyperparameters
C.2 Reward Setting
C.3 Does MAPPO work in Moving Out with sparse reward setting?
C.3.1 Dense Reward Setting
@@
Additionally, there are special rewards tailored for specific maps. In Map 11, for instance, two agents need to hold the two short sides of a rectangular item to more easily pass through a path successfully. Therefore, to encourage this, the reward calculation for the agents’ distance to this item has been modified: instead of being based on the distance to the item’s center point, it is now calculated based on the distance to its two short sides. This change is designed to encourage the agents to grasp the rectangle by its short ends.
@@ B. Event-based Rewards: Agent Holds an Item
Agent successfully holds an item Default reward for picking up $+0.5$
Exception: If another agent is already holding other middle or large item at this time $-0.5$ (total for this hold event)
Exception: If the item picked up was already located within a goal region $-0.5$ (total for this hold event)
@@ C. Event-based Rewards: Agent Unholds an Item
Agent successfully unholds an item Item is released inside a goal region $+0.5$
Item is released not inside a goal region $-0.5$
Exception: If another agent needs help, holding a large or middle item outside the goal region, at the moment of unholding. $+0.5$ (additive)
@@ D. Time-based Reward (Step Cost)
Each timestep Agent exists in the environment $-0.01$
Appendix D Comparative analysis of the Behaviors of BC and RL agents
@@
This distinction is evident in specific scenarios. For instance, on Map 5, both agents have their closest middle-sized items. However, human demonstration data frequently shows a strategy of first securing two smaller items before returning to move a middle-sized item together. A MAPPO agent, guided by a dense reward that incentivizes moving the nearest object, will typically prioritize the closer middle-sized item. If two such items are equidistant to respective agents (e.g., a pink agent targeting a yellow star and a blue agent targeting a blue circle), the initial actions will be independent. The coordination emerges when one agent successfully grasps a middle-sized item; the reward structure then incentivizes the other agent to assist with that specific item. Thus, the RL behavior can appear as a race to secure a primary middle-sized object, with the "loser" then being redirected by rewards to help the "winner." BC models on Map 5, however, reflect the diversity of the human dataset. This dataset contains instances of both "small-items-first" and "middle-item-first" strategies. Consequently, a BC agent might exhibit behaviors where one agent targets a middle-sized item while the other simultaneously attempts to move a small item, reflecting a momentary misalignment as different agents emulate distinct strategies observed in the human data.
Map 10 further illustrates these differences. Here, two agents might each have two items at an equal distance, making multiple initial moves potentially optimal. In our MAPPO training, agents often exhibit initial movements that appear somewhat exploratory or randomized until one agent commits to and grasps a large item. At this point, the dense reward system effectively directs the other agent to provide assistance. Conversely, BC models on Map 11 Four Corners tend to display more decisive and rapidly aligned behavior from the start. Observations of the human dataset for this map revealed a common leader-follower dynamic, where one player (e.g., the blue agent) consistently follows the lead of the other (e.g., the pink agent). If the pink agent, for example, decisively moves towards an upper pink square, the blue agent often follows suit immediately to assist. As a result, BC models rarely exhibit prolonged periods of uncoordinated or hesitant movement before aligning on a common goal.
@@
Appendix E Details of Evaluation Metrics
@@
where $w_{i}=1$ (small) or $2$ (middle/large). Range: [0,1].
@@
Appendix F Implementation Details
F.1 Environment Details
F.1.1 Observation Encoding
@@ F.1.2 Action Encoding
  • The grasping action: $1$ means grasp or release, 00 means no change.
F.1.3 Evaluation Setting
F.2 Baseline Details
  • Diffusion Policy: We follow the original implementation by [24] for the model architecture, which employs a 1D U-Net to generate action sequences. The observation, prediction, and executable horizons are set to 2, 8, and 4, respectively. Training is performed using the Adam optimizer with 1k epochs, 1024 batch size, and 0.001 learning rate. The diffusion steps are 36. The grasp action is encoded by one-hot encoding.
@@ F.3 BASS Details
In behavior augmentation, we add noise with a mean of 00 and a standard deviation of $0.002$. In recombination sub-trajectories, since two points in a continuous space are almost never the same, we set a tolerance value. We discretize the environment into a $48\times 48$ grid. If the robot’s start and end points are in the same grid cell, we treat them as the same point.
@@
Many maps have walls, so we cannot use Euclidean distance. To improve efficiency, we discretize the environment into a $48\times 48$ grid. We use the BFS algorithm to compute the distance from the item to the Goal Region.
@@
Appendix G Data Collection: Training Data
@@
For Task 1, which focuses on human behavior diversity, we recruited 36 participants, forming 18 groups of two. Before data collection, each group underwent a 10-minute practice session to familiarize itself with the environment. The remaining 50 minutes were dedicated to data collection. Each pair played each map three times, then switched agents and played three more times, resulting in six demonstrations per map. If a group completed all maps, they contributed a total of $12\times 6=72$ human demonstrations. However, not all groups completed the full set, with some collecting only 3 to 5 demonstrations per map. Additionally, we removed low-quality demonstrations where performance was significantly poor. In total, we collected 1,000 valid human demonstrations for this task.
For Task 2, which evaluates adaptation to physical constraints, we worked with four expert players who were highly familiar with the environment. Each map had randomized object properties, ensuring variation in shape, size, and mass. Each map was played 60 times, resulting in $60\times 12=720$ human demonstrations.
@@
Appendix H Computing Resources
H.1 Training
H.1.1 Behavior Cloning
@@ H.1.2 MAPPO
H.2 Inference Speed of DP/BASS
@@
Appendix I Human Study: Playing with Models
I.1 Human Study Procedure
@@ I.2 Questionnaire
Appendix J Results of Different Methods with BASS
@@
Appendix K Map Analysis
K.1 Coordination

### v2

@@ C.1 Hyperparameters
Appendix D Reward Setting
D.1 Dense Reward Setting
@@
Additionally, there are special rewards tailored for specific maps. In Map 11 (Four Corners), for instance, two agents need to hold the two short sides of a rectangular item to more easily pass through a path successfully. Therefore, to encourage this, the reward calculation for the agents’ distance to this item has been modified: instead of being based on the distance to the item’s center point, it is now calculated based on the distance to its two short sides. This change is designed to encourage the agents to grasp the rectangle by its short ends.
@@ B. Event-based Rewards: Agent Holds an Item
Agent successfully holds an item Default reward for picking up +0.5+0.5+ 0.5
Exception: If another agent is already holding other middle or large item at this time 0.5-0.5- 0.5 (total for this hold event)
Exception: If the item picked up was already located within a goal region 0.5-0.5- 0.5 (total for this hold event)
@@ C. Event-based Rewards: Agent Unholds an Item
Agent successfully unholds an item Item is released inside a goal region +0.5+0.5+ 0.5
Item is released not inside a goal region 0.5-0.5- 0.5
Exception: If another agent needs help, holding a large or middle item outside the goal region, at the moment of unholding. +0.5+0.5+ 0.5 (additive)
@@ D. Time-based Reward (Step Cost)
Each timestep Agent exists in the environment 0.01-0.01- 0.01
D.2 Does MAPPO work in Moving Out with sparse reward setting?
Appendix E Comparative analysis of the Behaviors of BC and RL agents
@@
This distinction is evident in specific scenarios. For instance, on Map 6 (Distance Priority), both agents have their closest middle-sized items. However, human demonstration data frequently shows a strategy of first securing two smaller items before returning to move a middle-sized item together. A MAPPO agent, guided by a dense reward that incentivizes moving the nearest object, will typically prioritize the closer middle-sized item. If two such items are equidistant to respective agents (e.g., a pink agent targeting a yellow star and a blue agent targeting a blue circle), the initial actions will be independent. The coordination emerges when one agent successfully grasps a middle-sized item; the reward structure then incentivizes the other agent to assist with that specific item. Thus, the RL behavior can appear as a race to secure a primary middle-sized object, with the "loser" then being redirected by rewards to help the "winner." BC models on Map 6 (Distance Priority), however, reflect the diversity of the human dataset. This dataset contains instances of both "small-items-first" and "middle-item-first" strategies. Consequently, a BC agent might exhibit behaviors where one agent targets a middle-sized item while the other simultaneously attempts to move a small item, reflecting a momentary misalignment as different agents emulate distinct strategies observed in the human data.
Map 11 (Four Corners) further illustrates these differences. Here, two agents might each have two items at an equal distance, making multiple initial moves potentially optimal. In our MAPPO training, agents often exhibit initial movements that appear somewhat exploratory or randomized until one agent commits to and grasps a large item. At this point, the dense reward system effectively directs the other agent to provide assistance. Conversely, BC models on Map 11 (Four Corners) tend to display more decisive and rapidly aligned behavior from the start. Observations of the human dataset for this map revealed a common leader-follower dynamic, where one player (e.g., the blue agent) consistently follows the lead of the other (e.g., the pink agent). If the pink agent, for example, decisively moves towards an upper pink square, the blue agent often follows suit immediately to assist. As a result, BC models rarely exhibit prolonged periods of uncoordinated or hesitant movement before aligning on a common goal.
@@
Appendix F Details of Evaluation Metrics
@@
where $w_{i}=1$ (small) or 222 (middle/large). Range: [0,1].
@@
Appendix G Implementation Details
G.1 Environment Details
G.1.1 Observation Encoding
@@ G.1.2 Action Encoding
  • The grasping action: 111 means grasp or release, 0 means no change.
G.2 Baseline Details
  • Diffusion Policy: We follow the original implementation by Chi et al. [2023] for the model architecture, which employs a 1D U-Net to generate action sequences. The observation, prediction, and executable horizons are set to 2, 8, and 4, respectively. Training is performed using the Adam optimizer with 1k epochs, 1024 batch size, and 0.001 learning rate. The diffusion steps are 36. The grasp action is encoded by one-hot encoding.
@@ G.3 BASS Details
In behavior augmentation, we add noise with a mean of 0 and a standard deviation of 0.0020.0020.002. In recombination sub-trajectories, since two points in a continuous space are almost never the same, we set a tolerance value. We discretize the environment into a $\times 48$ 48 grid. If the robot’s start and end points are in the same grid cell, we treat them as the same point.
@@
Many maps have walls, so we cannot use Euclidean distance. To improve efficiency, we discretize the environment into a $\times 48$ 48 grid. We use the BFS algorithm to compute the distance from the item to the Goal Region.
@@
Appendix H Data Collection: Training Data
@@
For Task 1, which focuses on human behavior diversity, we recruited 36 participants, forming 18 groups of two. Before data collection, each group underwent a 10-minute practice session to familiarize itself with the environment. The remaining 50 minutes were dedicated to data collection. Each pair played each map three times, then switched agents and played three more times, resulting in six demonstrations per map. If a group completed all maps, they contributed a total of 12×6=7212$\times$ 6=7212 × 6 = 72 human demonstrations. However, not all groups completed the full set, with some collecting only 3 to 5 demonstrations per map. Additionally, we removed low-quality demonstrations where performance was significantly poor. In total, we collected 1,000 valid human demonstrations for this task.
For Task 2, which evaluates adaptation to physical constraints, we worked with four expert players who were highly familiar with the environment. Each map had randomized object properties, ensuring variation in shape, size, and mass. Each map was played 60 times, resulting in 60×12=72060$\times$ 12=72060 × 12 = 720 human demonstrations.
@@
Appendix I Computing Resources
I.1 Training
I.1.1 Behavior Cloning
@@ I.1.2 MAPPO
I.2 Inference Speed of DP/BASS
@@
Appendix J Human Study: Playing with Models
J.1 Human Study Procedure
@@
Appendix K Results of Different Methods with BASS
@@
Appendix L Map Analysis
L.1 Coordination